What is the annual 1.5-year theoretical spot rate if the market implied six-month, 1-year and 1.5-year forward rates are 4.00%, 4.20%, and 5.00%?
Answer #1

Implied 1.5-year forward rates = [1.5*T1.5-0.5*T0.5]/1.5-0.5

0.05 = [1.5*T1.5-0.5*0.04]

0.05+0.02 = 1.5*T1.5

T1.5 = 0.07/1.5 = 0.04667 = 4.667%